1/z= (1/(3-6i)) + (1/(5+4i)) please show work thanks!!!
Answer by stanbon(75887)

1/z= (1/(3-6i)) + (1/(5+4i))
--------------------------
1/z = (3+6i)/(9+36) + (5-4i)/(25+16)
------
1/z = (1+2i)/15 + (5-4i)/41
-----
1/z = [41(1+2i) + 15(5-4i)]/(15*41)
-----
1/z = (116+22i)/(15*41)
---
z/1 = 56/(116+22i)
----
z = 56(116-22i)/(116^2+22^2)
----
z = [56(116-22i)/13940]
etc.
=============
Cheers,
Stan H.
--------------